OpenStudy (2bornot2b):

Prove that \[\huge{(e^w)^z=e^{[(w+2ni\pi)(z)]}}\] but \[\huge (e^w)^z\ne e^{(wz)}\] where w and z are complex.

OpenStudy (2bornot2b):

@shivam_bhalla

myininaya (myininaya):

\[\text{ Do you recall } e^{ i \theta}= \cos(\theta) +i \sin(\theta)\]

myininaya (myininaya):

\[e^{i \theta}=\cos(\theta)+i \sin(\theta)=\cos(\theta+2 n \pi)+ i \sin( \theta + 2 n \pi))=e^{i(\theta+2 n \pi)}\]

myininaya (myininaya):

And of course n is an integer

myininaya (myininaya):

Try to use that in this proof I would write w as a+bi by the way until the end when you write a+bi back as w

OpenStudy (2bornot2b):

So you want me to start with \(\large (e^{a+ib})^z\) Sorry for the late reply @myininaya I was away.

OpenStudy (2bornot2b):

What do I do next @myininaya ?

OpenStudy (kinggeorge):

The first thing you need to show is not too bad. Since \(w\) is complex, \(w=w+2ni\pi \). Hence\[\Large (e^w)^z=e^{(w+2ni\pi)~z}\]by using a simple substitution.

OpenStudy (2bornot2b):

@KingGeorge, is it just substitution? Perhaps you didn't notice I have replaced the ( )

OpenStudy (kinggeorge):

Herp derp.

OpenStudy (kinggeorge):

My ideas aren't panning out :(

OpenStudy (anonymous):

What do you mean by complex? \(w=x+iy\)?

OpenStudy (2bornot2b):

@Ishaan94, Yes.

OpenStudy (anonymous):

Does the fact that w and z are complex numbers restrict the use of euler's identity?

OpenStudy (experimentx):

why \( (e^w)^z \not = e^{wz} \) ???

OpenStudy (2bornot2b):

I don't think so @Ishaan94 @experimentX see bullet three after visiting the following link http://en.wikipedia.org/wiki/Exponentiation#Failure_of_power_and_logarithm_identities

OpenStudy (2bornot2b):

Certainly you can use euler's identity

OpenStudy (anonymous):

I know that \(e^{x} =e^{y}\) doesn't imply that x equals y, but I still don't understand it. Hmm okay maybe I know why. \(e^{i\theta} = \cos \theta +i\sin \theta\). \[\sin 30 = \sin150,\,\text{but}\, 150\neq30.\]\[\sin \theta=\sin (2n\pi + \theta)\]

OpenStudy (anonymous):

You can only have one unique value of theta in a period by adding 2npi you change the period. Or, you complete the circle and reach the point where you started from.

OpenStudy (2bornot2b):

But what does it have to do with this complex identity ?

OpenStudy (anonymous):

hmm, maybe you have to go back to the definition in terms of complex log

OpenStudy (2bornot2b):

@estudier Do you mean the definition of \(a^b\) where both a and b are complex?

OpenStudy (anonymous):

Problem seems to be you cannot equate a multi value f with a single value f....

OpenStudy (anonymous):

A lot. This is how you represent a complex number in complex plane. You can only get the same theta value if you complete circle.